# Attractors for Singular-Degenerate Porous Medium Type Equations Arising in Models for Biofilm Growth

### v1

@@ 3 Existence of Global Attractors
    we have $S(T)u_{0}=S(T-\xi)S(\xi)u_{0}\in\mathcal{K}$. By the definition of $\mathcal{K}$, we observe that $\mathcal{K}=Z([0,T]\times K)$, where $Z(t,y):=S(t)y$. By the relative compactness of $K$, it follows that $[0,T]\times K$ is also relatively compact. Furthermore, the function $Z$ is continuous in $L^{1}(\Omega)$. Indeed, for any sequences $\{y_{k}\}_{k=1}^{\infty}\subset L^{1}(\Omega)$ and $\{t_{k}\}_{k=1}^{\infty}\subset[0,\infty)$ such that $y_{k}\to y$ and $t_{k}\to t$ as $k\to\infty$, by (2.13), we get
@@
    The first term on the right-hand side converges to zero as $k\to\infty$ and the second term as well, due to the Lipschitz continuity of $\mathcal{S}$ with respect to $t$ in $L^{1}(\Omega)$, see Theorem 2.1. Consequently, the set $\mathcal{K}$ is relatively compact in $L^{1}(\Omega)$. Therefore, the set $\mathcal{S}(T)\mathcal{X}$ is relatively compact in $L^{1}(\Omega)$. Since $\mathcal{X}$ is positively invariant, we have
@@
  • Proof of Theorem 2.3:.
    By Theorem 2.1, the semigroup $\{\mathcal{S}(t)\}_{t\geq 0}$ is continuous in $L^{1}(\Omega)$ with respect to $t$. Therefore, using that $\mathcal{X}$ is a bounded absorbing set together with Lemma 3.1, we can apply the abstract theorem [33, Theorem 1.1] on the existence of global attractors to conclude the proof of the theorem. ∎
Finally, we present the proof of the existence of the global attractor to the coupled problem (2.16).
  • Proof of Theorem 2.4:.
    Analogously as in the proof of Theorem 2.3 we obtain the continuity of the semigroup $\{\mathcal{T}(t)\}_{t\geq 0}$ in $L^{1}(\Omega)\times L^{1}(\Omega)$ with respect to $t$ and the existence of a bounded absorbing set, which is $\mathcal{Y}$ by Theorem 2.2. Therefore, it remains to show the asymptotic compactness of the semigroup. We only present a sketch of the proof as the arguments are the same as in the proof of Lemma 3.1.
@@
    $||\phi(z_{1})||_{L^{\infty}(\Omega)}+||z_{2}||_{L^{\infty}(\Omega)}+||\nabla\phi(z_{1})||_{L^{2}(\Omega)}+||\nabla z_{2}||_{L^{2}(\Omega)}\leq R\},$
    for any $R>0$, is relatively compact in $L^{1}(\Omega)\times L^{1}(\Omega)$. Moreover, by Theorem 2.2 (see also Remark 2.5), for any $T>0$, there exists $\xi\in[0,T]$ such that $\mathcal{T}(\xi)(u_{0},v_{0})=(u(\xi),v(\xi))\in K$, $K=K(C)$ for some constant $C>0$. Hence, it follows that $\mathcal{T}(T)(u_{0},v_{0})\in\mathcal{K}$, for any $(u_{0},v_{0})\in\mathcal{Y}$, where the set
@@
    is relatively compact in $L^{1}(\Omega)\times L^{1}(\Omega)$. Therefore, as in the proof of Lemma 3.1 we conclude that the set $\bigcup\limits_{t\geq T}\mathcal{T}(t)\mathcal{Y}$ is relatively compact. The abstract theorem [33, Theorem 1.1] on the existence of the global attractor can be applied, which completes the proof. ∎
@@ 4 Existence of Exponential Attractors
We first consider the scalar problem (2.1) and show the existence of an exponential attractor for the corresponding semigroup $\{\mathcal{S}(t)\}_{t\geq 0}$, i.e. we prove Theorem 2.5. To this end, we follow and adapt the arguments in [11, Theorems 2.1 and 3.1] to our case. However, different from [11], where first the finite fractal dimension of the global attractor was shown and then the exponential attractor was constructed from scratch, we apply an abstract existence result for exponential attractors for discrete time semigroups. Subsequently, the exponential attractor for the continuous time semigroup can be constructed in a standard manner, see [5, 11, 31]. This significantly shortens the existence proof for the exponential attractor, and the finite fractal dimension of the global attractor is an immediate consequence. More specifically, the existence result for discrete time exponential attractors we will use is the following proposition which is a slight generalization of [5, Theorem 2.1], see also [11, Remark 3.2]. Here, we give a specific estimate of the fractal dimension of the exponential attractor in terms of the constants in the relevant estimates.
@@
All estimates obtained from now on can be justified by approximating the solutions of (2.1) and (2.16) by smooth solutions of non-degenerate problems.
@@
$\mathcal{C}:=\overline{\bigcup_{t\geq 0}S(t)\mathcal{X}},$
where the closure is taken in $C^{\alpha}(\bar{\Omega})$. By (2.20), we have $\lVert\mathcal{C}\rVert_{C^{\alpha}(\bar{\Omega})}\leq M$ and due to the construction of $\mathcal{C}$, we have $S(t)\mathcal{C}\subset\mathcal{C}$ for all $t\geq 0$. Let $u_{0}\in\mathcal{C},u_{0}\neq 0$ be arbitrary. We introduce the sets
@@
    Recall that by (2.20), the solution $u(t):=\mathcal{S}(t)u_{0}$ belongs to $C^{\alpha}([0,\infty)\times\overline{\Omega})$. Hence, for any $T>0$ we have
@@
    Then, by (4.7)2italic-(4.7subscriptitalic-)$_{2}$, it follows that
    $|v(t,x)|\geq|u(t,x)|-|v(t,x)-u(t,x)|>\theta/2-C_{0}\varepsilon^{\gamma}\quad\text{for}\ x\in L(\theta),\ t\in[0,T].$
    Hence, for all sufficiently small $\varepsilon$ independent of $u_{0}\in\mathcal{C}$, considering (4.7)1italic-(4.7subscriptitalic-)$_{1}$, we obtain (4.3). ∎
Next, we show a smoothing property for the difference of two solutions of (2.1). We first consider the domain $L(\theta)$, where the equation (2.1)1italic-(2.1subscriptitalic-)$_{1}$ is non-degenerate. The following lemma is analogous to the standard interior regularity estimates for linear parabolic problems.
@@
    where $W^{(1,2),r}(D)$ denotes the anisotropic Sobolev space consisting of functions whose $t$-derivatives up to order $1$ and $x$-derivatives up to order $2$ belong to $L^{r}(D)$. The constant $C_{r,t_{0}}$ depends on $r$ and $t_{0}$, but is independent of $u_{0},\varepsilon$ and the trajectories $u_{1}$ and $u_{2}$. It suffices to verify (4.10) for $u=u_{1}$ since the proof is analogous for $u=u_{2}$. To this end, we introduce $v(t,x):=\phi(u(t,x))$. Then, considering (4.3)2italic-(4.3subscriptitalic-)$_{2}$ and that $\phi(u)$ is non-degenerate if $|u|>\theta/4>0$, the function $v$ solves the equation
@@
    where $a(t,x):=\phi^{\prime}(u(t,x))$. Furthermore, by (2.20), the continuity of $\phi^{\prime}$ and $f$, the function $a$ is uniformly Hölder continuous with respect to $t$ and $x$ and the function $af(\cdot,u)$ is uniformly bounded in $L^{\infty}(\Omega_{T})$. Additionally, due to (2.6) and (4.3)2italic-(4.3subscriptitalic-)$_{2}$, we have
@@
    where the constant $C$ is also independent of $u_{0}$ and $u$. Hence, we can use an $L^{r}$-interior regularity estimate for the solution of the linear non-degenerate equation (4.11) (see [11, Proposition A.4]), and obtain, by (4.2), for any $r>2$ and $t_{0}\in(0,T)$,
@@
    $\quad\lVert u\rVert_{W^{(1,2),r}((t_{0}/2,T)\times L(2\theta))}^{r}$ $=\lVert\phi^{-1}(v)\rVert_{W^{(1,2),r}((t_{0}/2,T)\times L(2\theta))}^{r}$
    $=\lVert\phi^{-1}(v)\rVert_{L^{r}((t_{0}/2,T)\times L(2\theta))}^{r}+\left\|\frac{v_{t}}{\phi^{\prime}(u)}\right\|_{L^{r}((t_{0}/2,T)\times L(2\theta))}^{r}$
    $+\sum_{i=1}^{n}\left\|\frac{v_{x_{i}}}{\phi^{\prime}(u)}\right\|_{L^{r}((t_{0}/2,T)\times L(2\theta))}^{r}$
@@
    To prove estimate (4.9) we consider $w=u_{1}-u_{2}$ in $[t_{0}/2,T]\times L(2\theta)$ which satisfies
@@
    $||l_{1}||_{C^{\alpha}([0,T]\times\bar{\Omega})}\leq C,$ (4.16)
    where $C$ is independent of $u_{1}$ and $u_{2}$. Furthermore, by (4.10), we have
    $||\partial_{t}l_{1}||_{L^{r}((t_{0}/2,T)\times L(2\theta))}\leq C\sum_{i=1}^{2}||\partial_{t}u_{i}||_{L^{r}((t_{0}/2,T)\times L(2\theta))}\leq C_{r,t_{0}}^{\prime\prime\prime},$ (4.17)
    for some constant $C_{r}^{\prime\prime\prime}>0$ independent of $u_{0},u_{1}$ and $u_{2}$. Also, by (4.3)2italic-(4.3subscriptitalic-)$_{2}$ and (2.6), we have
@@
    $||Z||_{W^{(1,2),q}((t_{0},T)\times L(3\theta))}\leq C_{q,t_{0}}||Z||_{L^{1}((t_{0}/2,T)\times L(2\theta))}\leq C_{q,t_{0}}^{\prime}||w||_{L^{1}(\Omega_{T})},$ (4.20)
@@
    $||w||_{C^{\alpha}([t_{0},T]\times\overline{L(3\theta)})}=\left\|\frac{Z}{l_{1}}\right\|_{C^{\alpha}([t_{0},T]\times\overline{L(3\theta)})}$ $\leq C||Z||_{C^{\alpha}([t_{0},T]\times\overline{L(3\theta)})}$
    $\leq C||Z||_{W^{(1,2),q}((t_{0},T)\times L(3\theta))}$
    $\leq C_{q,t_{0}}||w||_{L^{1}(\Omega_{T})}\leq C_{q,t_{0}}e^{LT}||w(0)||_{L^{1}(\Omega)},$
@@
$||u_{1}(T)-u_{2}(T)||_{L^{1}(S(4\theta))}\leq\frac{1-\delta}{2}||u_{1}(0)-u_{2}(0)||_{L^{1}(\Omega)}+C||u_{1}-u_{2}||_{L^{1}([t^{*},T]\times L(4\theta))}$ (4.21)
@@
    $w_{t},\psi\text{sgn}(w)\right>=\left<\Delta(l_{1}(t)w),\psi\text{sgn}(\psi l_{1}(t)w)\right>+\left<l_{2}(t)w,\psi sgn(w)\right>.$
@@
    $\partial_{t}\left<\psi,|w|\right>$ =$\Delta(\psi l_{1}(t)w),\text{sgn}(\psi l_{1}(t)w)\right>-2\left<\nabla\psi\cdot\nabla(l_{1}(t)w),\text{sgn}(\psi l_{1}(t)w)\right>$
    $l_{1}(t)w\Delta\psi,\text{sgn}(\psi l_{1}(t)w)\right>+\left<l_{2}(t)w,\psi\text{sgn}(w)\right>.$
    Then, applying Kato’s inequality $(\Delta v,\text{sgn}v)\leq 0$ and integration by parts to the second term on the right hand side, we deduce that
@@
    Therefore, using (4.3)1italic-(4.3subscriptitalic-)$_{1}$, (4.6) and (4.16), and the fact that $\Delta\psi(x)=0$ for $x\in S(4\theta)$, we conclude that
@@
    $\leq\frac{1-\delta}{2}\lVert u_{1}(0)-u_{2}(0)\rVert_{L^{1}(\Omega)}+C||u_{1}-u_{2}||_{L^{1}([t_{*},T]\times L(4\theta))}$
@@
    is compact and uniform in the sense of the Kolmogorov $\varepsilon$-entropy. Hence, we can apply Proposition 4.1 to the discrete time semigroup $S(n)=\mathcal{S}(Tn),n\in\mathbb{N}_{0}$, with $E=L^{1}(\Omega),$ $K=\mathcal{C},u_{*}=u_{0},$ $X_{u_{*},\varepsilon}=C^{\alpha}([t_{*},T]\times\overline{L(3\theta)})$ and $Y_{u_{*},\varepsilon}=C([t_{*},T]\times\overline{L(7\theta/2)})$. Moreover, the mapping $\mathcal{T}_{u_{*},\varepsilon}$ is defined by $\mathcal{T}_{u_{*},\varepsilon}(v_{0})=v|_{[t_{*},T]\times\overline{L(3\theta)}}$, i.e. the restriction of the solution $v$ with initial data $v_{0}\in B^{L^{1}(\Omega)}_{\varepsilon}(u_{0})\cap\mathcal{C}$ to the set $[t_{*},T]\times\overline{L(3\theta)}.$ Assumption (2.20) and the estimates (4.23) and (4.24) show that the hypotheses of Proposition 4.1 hold with the constants $\mu,\kappa$ and $\eta=\frac{1-\delta}{2}<\frac{1}{2}$. Hence, for any $\nu\in(0,\frac{\delta}{2})$ the discrete time semigroup $\{S(n)\}_{n\in\mathbb{N}_{0}}$ possesses an exponential attractor $\mathcal{M}_{\nu}^{d}\subset\mathcal{C}$ and its fractal dimension is bounded by
@@
    As seen in the proof of [11, Theorem 3.1] and [5, Theorem 2.1], the discrete exponential attractor is the closure in $L^{1}(\Omega)$ of a countable set in $\mathcal{C}$, $\mathcal{M}_{\nu}^{d}=\overline{\widetilde{\mathcal{M}}_{\nu}^{d}}$, where ~$\widetilde{\mathcal{M}}_{\nu}^{d}=\bigcup_{n\in\mathbb{N}_{0}}V^{n}$ with $V^{n}\in\mathcal{C}.$
    Finally, to obtain an exponential attractor for the continuous time semigroup we follow the construction in [31]. We set $\mathcal{M}_{\nu}=\overline{\widetilde{\mathcal{M}}_{\nu}}$, where the closure is taken in $L^{1}(\Omega)$ and
    ~$\widetilde{\mathcal{M}}_{\nu}=\bigcup_{t\in[0,T]}\mathcal{S}(t)\widetilde{\mathcal{M}}_{\nu}^{d}.$
@@
    see [31, Theorem 3.4]. This provides the estimate for the fractal dimension of the exponential attractor in Theorem 2.5 with $J=[t_{*},T],$ $\Omega_{1}=L(3\theta),\Omega_{2}=L(7\theta/2)$, $\eta_{1}=1-\delta+2\nu$ and $\eta_{2}=\frac{\nu}{2\kappa},$ which completes the proof. ∎
@@
    By (2.23), the solution $w=(u,v)$ belongs to $C^{\alpha}$ with respect to $t$ and $x$. Then, in the same manner as in Lemma 4.2, we conclude that for an arbitrary solution $\tilde{w}:=(\tilde{u},\tilde{v})$ of the problem (2.16) with initial data $\tilde{w}_{0}=(\tilde{u}_{0},\tilde{v}_{0})\in B^{L^{1}(\Omega)\times L^{1}(\Omega)}_{\varepsilon}(w_{0})\cap\mathcal{C}$, the following estimates hold, uniformly with respect to $\tilde{w}_{0}$,
@@
    where $\psi$ is the cut-off function defined in (4.5), and integrate over $\Omega$. Then, by Kato’s inequality, (4.25)1italic-(4.25subscriptitalic-)$_{1}$, (2.17) and (4.6), we obtain
@@
    for all solutions $w_{1},w_{2}$ with initial data $w_{i}(0)\in B^{L^{1}(\Omega)\times L^{1}(\Omega)}_{\varepsilon}(w_{0})\cap\mathcal{C}$, $i=1,2,$ where $10<\delta<$ 1, $T>0$, $\kappa=C_{t_{*}}$ and $\mu$ are independent of $\varepsilon\leq\varepsilon_{0}$, $w_{0}$ and $w_{1},w_{2}$. Hence, we can apply Proposition 4.1 to the discrete time semigroup $\Pi(n)=\mathcal{T}(Tn),n\in\mathbb{N}_{0}$, with $E=L^{1}(\Omega)\times L^{1}(\Omega),$ $K=\mathcal{C},u_{*}=w_{0},$ $X_{u_{*},\varepsilon}=C^{\alpha}([t_{*},T]\times\overline{L(3\theta)})\times C^{\alpha}([t_{*},T]\times\bar{\Omega}))$ and $Y_{u_{*},\varepsilon}=C([t_{*},T]\times\overline{L(7\theta/2)})\times C([t_{*},T]\times\bar{\Omega})$. Moreover, the mapping $\mathcal{T}_{u_{*},\varepsilon}$ is defined by $\mathcal{T}_{u_{*},\varepsilon}(\tilde{w}_{0})=(\tilde{u}|_{[t_{*},T]\times\overline{L(3\theta)}},\tilde{v}|_{[t_{*},T]\times\overline{\Omega}})$, i.e. the restriction of the solution $\bar{w}=(\bar{u},\bar{v})$ with initial data $\bar{w}_{0}\in B^{L^{1}(\Omega)\times L^{1}(\Omega)}_{\varepsilon}(w_{0})\cap\mathcal{C}$ to the sets $[t_{*},T]\times\overline{L(3\theta)}$ and $[t_{*},T]\times\Omega$, respectively. Assumption (2.23) and the estimates (4.23) and (4.24) show that the hypotheses of Proposition 4.1 hold with $\mu,\kappa$ and $\eta=\frac{1-\delta}{2}<\frac{1}{2}$. Hence, for any $\nu\in(0,\frac{\delta}{2})$ the discrete time semigroup $\{\Pi(n)\}_{n\in\mathbb{N}_{0}}$ possesses an exponential attractor $\mathcal{M}_{\nu}^{d}\subset\mathcal{C}$ and its fractal dimension is bounded by

### v2

@@ 3 Existence of Global Attractors
we have $\mathcal{S}(T)u_{0}=\mathcal{S}(T-\xi)\mathcal{S}(\xi)u_{0}\in\mathcal{K}$. By the definition of $\mathcal{K}$, we observe that $\mathcal{K}=Z([0,T]\times K)$, where $Z(t,y):=\mathcal{S}(t)y$. By the relative compactness of $K$, it follows that $0,T]\times K$ is also relatively compact. Furthermore, the function $Z$ is continuous in $L^{1}(\Omega)$. Indeed, for any sequences $\{y_{k}\}_{k=1}^{\infty}\subset L^{1}(\Omega)$ and $\{t_{k}\}_{k=1}^{\infty}\subset[0,T]$ such that $y_{k}\to y$ and $t_{k}\to t$ as $k\to\infty$, by (2.11), we get
@@
The first term on the right-hand side converges to zero as $k\to\infty$ and the second term as well, due to the Lipschitz continuity of $\mathcal{S}$ with respect to $t$ in $L^{1}(\Omega)$, see Theorem 2.1. Consequently, the set $\mathcal{K}$ is relatively compact in $L^{1}(\Omega)$. Therefore, $u(T)=\mathcal{S}(T)u_{0}$ has a convergent subsequence which converges to $u$. Hence, considering that $C$ is independent of $u_{0}$, the set $\mathcal{S}(T)\mathcal{X}$ is relatively compact in $L^{1}(\Omega)$. Since $\mathcal{X}$ is positively invariant, we have
@@
By Theorem 2.1, the semigroup $\{\mathcal{S}(t)\}_{t\geq 0}$ is continuous in $L^{1}(\Omega)$ with respect to $t$. Therefore, using that $\mathcal{X}$ is a bounded absorbing set together with Lemma 3.6, we can apply the abstract theorem [7, Theorem 1.1] on the existence of global attractors to conclude the proof of the theorem. ∎
Finally, we present a sketch of the proof of the existence of the global attractor to the coupled problem (2.12). As for the scalar problem (2.1), we can consider smooth non-degenerate approximations and derive uniform estimates for the approximate solutions. The difficulties lie in the estimates for the degenerate equation which we provided in this section, the arguments for the semilinear equation are standard, see [21]. The semigroup in the phase space $\mathcal{Y}$ is then obtained analogously as in the scalar case, see also [21]. To prove the asymptotically compactness of the semigroup we use arguments from [24].
Analogously as in the proof of Theorem 2.3, we obtain the continuity of the semigroup $\{\mathcal{T}(t)\}_{t\geq 0}$ in $L^{1}(\Omega)\times L^{1}(\Omega)$ with respect to $t$ and the existence of a bounded absorbing set, which is $\mathcal{Y}$ by Theorem 2.2. Therefore, it remains to show the asymptotic compactness of the semigroup. We only present a sketch of the proof as the arguments are the same as in the proof of Lemma 3.6.
@@
for any $R>0$, is relatively compact in $L^{1}(\Omega)\times L^{1}(\Omega)$. Moreover, by an analogous estimate as (3.19) from Theorem 3.5 for the scalar problem, for any $T>0$, there exists $\xi\in[0,T]$ such that $\mathcal{T}(\xi)(u_{0},v_{0})=(u(\xi),v(\xi))\in K$, where $K=K(C)$ for some constant $C>0$, which is independent of $(u_{0},v_{0})$. Hence, it follows that $\mathcal{T}(T)(u_{0},v_{0})\in\mathcal{K}$, for any $(u_{0},v_{0})\in\mathcal{Y}$, where the set
@@
is relatively compact in $L^{1}(\Omega)\times L^{1}(\Omega)$. Therefore, as in the proof of Lemma 3.6 we conclude that the set $\bigcup\limits_{t\geq T}\mathcal{T}(t)\mathcal{Y}$ is relatively compact. The abstract theorem [7, Theorem 1.1] on the existence of the global attractor can be applied, which completes the proof. ∎
@@ 4 Existence of Exponential Attractors
We first consider the scalar problem (2.1) and show the existence of an exponential attractor for the corresponding semigroup $\{\mathcal{S}(t)\}_{t\geq 0}$, i.e. we prove Theorem 2.5. To this end, we follow and adapt the arguments in [19, Theorems 2.1 and 3.1] to our case. However, different from [19], where first the finite fractal dimension of the global attractor was shown and then the exponential attractor was constructed from scratch, we apply an abstract existence result for exponential attractors for discrete time semigroups. Subsequently, the exponential attractor for the continuous time semigroup can be constructed in a standard manner, see [18, 19, 20, 34]. This significantly shortens the existence proof for the exponential attractor, and the finite fractal dimension of the global attractor is an immediate consequence. More specifically, the existence result for discrete time exponential attractors we will use is the following proposition which is a slight generalization of [18, Theorem 2.1], see also [19, Remark 3.2]. Here, we give a specific estimate of the fractal dimension of the exponential attractor in terms of the constants in the relevant estimates.
@@
All estimates obtained from now on can be justified by approximating the solutions of (2.1) and (2.12) by smooth solutions of non-degenerate problems, see Section 3.
@@
$\mathcal{C}:=\overline{\bigcup_{t\geq 1}\mathcal{S}(t)\mathcal{X}},$
where the closure is taken in $C^{\alpha}(\bar{\Omega})$. By (2.17), we have $\lVert\mathcal{C}\rVert_{C^{\alpha}(\bar{\Omega})}\leq M$ and due to the construction of $\mathcal{C}$, we have $\mathcal{S}(t)\mathcal{C}\subset\mathcal{C}$ for all $t\geq 0$. Let $u_{0}\in\mathcal{C},u_{0}\neq 0$ be arbitrary. We introduce the sets
@@
Recall that by (2.17), the solution $u(t):=\mathcal{S}(t)u_{0}$ belongs to $C^{\alpha}([0,T]\times\overline{\Omega})$. Hence, for any $T>0$ we have
@@
[…]
$|v(t,x)|\geq|u(t,x)|-|v(t,x)-u(t,x)|>\theta/2-C\varepsilon^{\gamma}\quad\text{for}\ x\in L(\theta),\ t\in[0,T].$
[…]
@@
[…]
@@
[…]
@@
where the constant $C$ is also independent of $u_{0}$ and $u$. Hence, we can use an $L^{r}$-interior regularity estimate for the solution of the linear non-degenerate equation (4.11) (see [19, Proposition A.4]), and obtain, by (4.2), for any $r>2$ and $t_{0}\in(0,T)$,
@@
$\lVert u\rVert_{W^{(1,2),r}((t_{0}/2,T)\times L(2\theta))}^{r}=\lVert\phi^{-1}(v)\rVert_{W^{(1,2),r}((t_{0}/2,T)\times L(2\theta))}^{r}$
$=\lVert\phi^{-1}(v)\rVert_{L^{r}((t_{0}/2,T)\times L(2\theta))}^{r}+\left\|\frac{v_{t}}{\phi^{\prime}(u)}\right\|_{L^{r}((t_{0}/2,T)\times L(2\theta))}^{r}+\sum_{i=1}^{n}\left\|\frac{v_{x_{i}}}{\phi^{\prime}(u)}\right\|_{L^{r}((t_{0}/2,T)\times L(2\theta))}^{r}$
@@
To prove estimate (4.9), we consider $w=u_{1}-u_{2}$ in $t_{0}/2,T]\times L(2\theta)$ which satisfies
@@
$\begin{split}l_{1}(t)&:=\int_{0}^{1}\phi^{\prime}(su_{1}(t)+(1-s)u_{2}(t))ds,\\ l_{2}(t)&:=\int_{0}^{1}f_{u}(\cdot,su_{1}(t)+(1-s)u_{2}(t))ds.\end{split}$ (4.15)
$l_{1}||_{C^{\alpha}([0,T]\times\bar{\Omega})}\leq C,$ (4.16)
where $C$ is independent of $u_{1}$ and $u_{2}$. Furthermore, by (4.10), we have
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
$\left<w_{t},\psi\text{sgn}(w)\right>=\left<\Delta(l_{1}(t)w),\psi\text{sgn}(\psi l_{1}(t)w)\right>+\left<l_{2}(t)w,\psi sgn(w)\right$>.
@@
$\partial_{t}\left<\psi,|w|\right$> $=\left<\Delta(\psi l_{1}(t)w),\text{sgn}(\psi l_{1}(t)w)\right>-2\left<\nabla\psi\cdot\nabla(l_{1}(t)w),\text{sgn}(\psi l_{1}(t)w)\right$>
$\quad-\left<l_{1}(t)w\Delta\psi,\text{sgn}(\psi l_{1}(t)w)\right>+\left<l_{2}(t)w,\psi\text{sgn}(w)\right$>.
Then, using the inequality $\langle\Delta v,\text{sgn}(v)\rangle\leq 0$ we obtain
$\partial_{t}\left<\psi,|w|\right>\leq\left<\Delta\psi,l_{1}(t)|w|\right>+\left<\psi,l_{2}(t)|w|\right>+2\left<l_{1}(t)w\nabla\psi,\nabla(\text{sgn}(w))\right$>,
@@
$\partial_{t}\left<\psi,|w(t)|\right>+\beta\left<\psi,|w(t)|\right$>
@@
[…]
@@
is compact and uniform in the sense of the Kolmogorov $\varepsilon$-entropy. Hence, we can apply Proposition 4.1 to the discrete time semigroup $S(n)=\mathcal{S}(Tn),n\in\mathbb{N}_{0}$, with $E=L^{1}(\Omega),$ $K=\mathcal{C},u_{*}=u_{0},$ $X_{u_{*},\varepsilon}=C^{\alpha}([t_{*},T]\times\overline{L(3\theta)})$ and $Y_{u_{*},\varepsilon}=C([t_{*},T]\times\overline{L(7\theta/2)})$. Moreover, the mapping $\mathcal{T}_{u_{*},\varepsilon}$ is defined by $\mathcal{T}_{u_{*},\varepsilon}(v_{0})=v|_{[t_{*},T]\times\overline{L(3\theta)}}$, i.e. the restriction of the solution $v$ with initial data $v_{0}\in B^{L^{1}(\Omega)}_{\varepsilon}(u_{0})\cap\mathcal{C}$ to the set $t_{*},T]\times\overline{L(3\theta)}.$ Assumption (2.17) and the estimates (4.23) and (4.24) show that the hypotheses of Proposition 4.1 hold with the constants $\mu,\kappa$ and $\eta=\frac{1-\delta}{2}<\frac{1}{2}$. Hence, for any $\nu\in(0,\frac{\delta}{2})$ the discrete time semigroup $\{S(n)\}_{n\in\mathbb{N}_{0}}$ possesses an exponential attractor $\mathcal{M}_{\nu}^{d}\subset\mathcal{C}$ and its fractal dimension is bounded by
@@
As seen in the proof of [19, Theorem 3.1] and [18, Theorem 2.1], the discrete exponential attractor is the closure in $L^{1}(\Omega)$ of a countable set in $\mathcal{C}$, $\mathcal{M}_{\nu}^{d}=\overline{\widetilde{\mathcal{M}}_{\nu}^{d}}$, where $\widetilde{\mathcal{M}}_{\nu}^{d}=\bigcup_{n\in\mathbb{N}_{0}}V^{n}$ with $V^{n}\in\mathcal{C}.$
Finally, to obtain an exponential attractor for the continuous time semigroup we follow the construction in [34], see also [20]. We set $\mathcal{M}_{\nu}=\overline{\widetilde{\mathcal{M}}_{\nu}}$, where the closure is taken in $L^{1}(\Omega)$ and
$\widetilde{\mathcal{M}}_{\nu}=\bigcup_{t\in[0,T]}\mathcal{S}(t)\widetilde{\mathcal{M}}_{\nu}^{d}.$
@@
see [34, Theorem 3.4]. This provides the estimate for the fractal dimension of the exponential attractor in Theorem 2.5 with $J=[t_{*},T],$ $\Omega_{1}=L(3\theta),\Omega_{2}=L(7\theta/2)$, $\eta_{1}=1-\delta+2\nu$ and $\eta_{2}=\frac{\nu}{2\kappa},$ which completes the proof. ∎
@@
[…]
@@
$\partial_{t}\left<\psi,|\xi_{1}(t)|\right>+\beta\left<\psi,|\xi_{1}(t)|\right$> $\leq\int\limits_{L(4\theta)}\Delta\psi(x)l_{1}(t,x)|\xi_{1}(t,x)|dx$
$\quad+\int\limits_{\Omega}|f(x,u_{2}(t,x),v_{1}(t,x))-f(x,u_{2}(t,x),v_{2}(t,x))||\psi(x)|dx$
@@
for all solutions $w_{1},w_{2}$ with initial data $w_{i}(0)\in B^{L^{1}(\Omega)\times L^{1}(\Omega)}_{\varepsilon}(w_{0})\cap\mathcal{C}$, $i=1,2,$ where $0<\delta<1$, $T>0$, $\kappa=C_{t_{*}}$ and $\mu$ are independent of $\varepsilon\leq\varepsilon_{0}$, $w_{0}$ and $w_{1},w_{2}$. Hence, we can apply Proposition 4.1 to the discrete time semigroup $\Pi(n)=\mathcal{T}(Tn),n\in\mathbb{N}_{0}$, with $E=L^{1}(\Omega)\times L^{1}(\Omega),$ $K=\mathcal{C},u_{*}=w_{0},$ $X_{u_{*},\varepsilon}=C^{\alpha}([t_{*},T]\times\overline{L(3\theta)})\times C^{\alpha}([t_{*},T]\times\bar{\Omega}))$ and $Y_{u_{*},\varepsilon}=C([t_{*},T]\times\overline{L(7\theta/2)})\times C([t_{*},T]\times\bar{\Omega})$. Moreover, the mapping $\mathcal{T}_{u_{*},\varepsilon}$ is defined by $\mathcal{T}_{u_{*},\varepsilon}(\tilde{w}_{0})=(\tilde{u}|_{[t_{*},T]\times\overline{L(3\theta)}},\tilde{v}|_{[t_{*},T]\times\overline{\Omega}})$, i.e. the restriction of the solution $\bar{w}=(\bar{u},\bar{v})$ with initial data $\bar{w}_{0}\in B^{L^{1}(\Omega)\times L^{1}(\Omega)}_{\varepsilon}(w_{0})\cap\mathcal{C}$ to the sets $t_{*},T]\times\overline{L(3\theta)}$ and $t_{*},T]\times\Omega$, respectively. Assumption (2.20) and the estimates (4.23) and (4.24) show that the hypotheses of Proposition 4.1 hold with $\mu,\kappa$ and $\eta=\frac{1-\delta}{2}<\frac{1}{2}$. Hence, for any $\nu\in(0,\frac{\delta}{2})$ the discrete time semigroup $\{\Pi(n)\}_{n\in\mathbb{N}_{0}}$ possesses an exponential attractor $\mathcal{M}_{\nu}^{d}\subset\mathcal{C}$ and its fractal dimension is bounded by